2. One of the features of money is its store of value. However, most people do not hold their wealth as currency. Given that currency is the most liquid type of asset, why don’t people hold all their wealth as currency?

3. Suppose the required reserve ratio is 20%. What is the maximum amount of total money supply that can be created from an initial deposit of $200? In general, why might the actual amount of total money creation be less than the maximum?

4. What is the change in the money supply when the Fed purchases $700 worth of bonds and the required reserve ratio is 14 percent assuming banks hold no excess reserves?

5. What is the change in the money supply when the Fed sold $700 worth of bonds and the required reserve ratio is 14 percent assuming banks hold no excess reserves?
Answer #1

2) People don't hold all their wealth as currency because they can earn interest rate on their idle money. People hold money to meet unforeseen events. If people hold all of their wealth as currency then opportunity cost is the interest rate which is foregone.

3) Required reserve = 20% of 200 = 40

Money left to make loans = Deposits - Required reserve = 200 - 40 = 160

Amount of total money supply created = 160 x 1/Required reserve ratio = 160 x 1/0.20 = 160 x 5 = 800

4) Change in money supply = Purchase of bonds x 1/Required reserve ratio = 700 x 1/0.14 = 5000(Increase in money supply)

5) Change in money supply = Sale of bonds x 1/Required reserve ratio = 700 x 1/0.14 = 5000 (Decrease in Money supply)